OpenStudy (anonymous):

**Help plzzz I give medals** 3x+4y=18 x+3y=11 Make a table of values for each linear equation. Include at least three rows of the table. Be sure to label each table.

OpenStudy (anonymous):

Before we can sub numbers in, rearrange the equations to be of the form y = ....

OpenStudy (anonymous):

is it 3

OpenStudy (anonymous):

No, I'll do the first one for you: 3x+4y=18 4y=18-3x y=(18-3x)/4 Do the same for the second equation then we can get going with the table.

OpenStudy (anonymous):

what did u do, u just put the number and didn't tell me what u did

OpenStudy (anonymous):

@tom982

OpenStudy (anonymous):

No, I rearranged the first equation to make it of the form y=....

OpenStudy (anonymous):

Do the same for the second equation

OpenStudy (anonymous):

dont know how to

OpenStudy (anonymous):

Don't know how to, or don't want to? This is very basic algebra.

OpenStudy (anonymous):

dont know

OpenStudy (anonymous):

y=9/2=3x/4

OpenStudy (anonymous):

@tom982

OpenStudy (anonymous):

Okay, let's work through my example for the first equation: 3x+4y=18 (initial equation) 4y=18-3x (subtract 3x from both sides) y=(18-3x)/4 (divide both sides by 4) We've gone from a horrible equation, to a nice one of the form y=... Now do the same for x+3y=11

OpenStudy (anonymous):

y=11/3-1/3x

OpenStudy (anonymous):

@tom982

OpenStudy (anonymous):

Yes!

OpenStudy (anonymous):

Okay, so now we have our two equations: y=(18-3x)/4 y=(11-x)/3 Let's plug in values of x and see what values of y come out: